OpenStudy (anonymous):

cathyg: is there a way to solve this with Algebra 2^x=x+3 I know I can set both sides =y and do 2nd calc intersect on a graph calc. …need Alge

OpenStudy (amistre64):

logarithms are algebra

OpenStudy (anonymous):

Unfortunately, there is no known method for finding exact solutions to such an equation as this. The vast majority of equations, in fact, cannot be solved analytically.

OpenStudy (amistre64):

log(2^x)=log(x+3) x log(2) = log(x+3) log(2) = log(xrt(x+3)) 2 = xrt(x+3) .... that doesnt help does it :)

OpenStudy (anonymous):

OK. I just wasn't sure abtrehearn. amistre64- I tried that!

OpenStudy (amistre64):

graph it as: 2^x -(x+3) = 0

OpenStudy (anonymous):

did that ,got 2 answ. just wanted to know about alge

OpenStudy (anonymous):

Thanks for your help. Is there a way to know if an equaiton cant be solvedd with algebra?

OpenStudy (amistre64):

the wolfram uses some complicated functions; so unless you know alot of the higher maths, you might just be stuck

OpenStudy (amistre64):

if your interested, check it out :) http://www.wolframalpha.com/input/?i=2^x+-%28x%2B3%29

OpenStudy (anonymous):

OK good deal. just stick w/ the graphing approach but may check out website too